# Robust Pronoun Use Fidelity with English LLMs: Are they Reasoning, Repeating, or Just Biased?

### v1

@@ Acknowledgements
The authors thank Timm Dill for several rounds of patient annotation, Aaron Mueller for early feedback on this work, and Marius Mosbach and Vlad Niculae for discussions about math, plots and framing. Vagrant Gautam received funding from the BMBF’s (German Federal Ministry of Education and Research) SLIK project under the grant 01IS22015C. The work of Anne Lauscher is funded under the Excellence Strategy of the German Federal Government and States.
Appendix A List of occupations
@@
Appendix B Context Template Construction
For each case, we create 10 explicit templates which explicitly demonstrate the coreference between an individual and a pronoun using a subordinate clause, and 10 implicit templates, simple sentences which only contain a pronoun as the subject. The first two sentences of the context (i.e., the introduction and the first distractor) are always sampled from the explicit templates, and the rest are sampled from the implicit templates.
@@ Appendix B Context Template Construction
After making a choice for the first distractor template, we fill it with any of the three remaining pronouns and then we remove this template’s index from our pool, but re-add the index of the introductory template. This is because subsequent distractor templates always use implicit templates. For example, if we chose $exp\_neg_{4}$ as our first distractor template, we would now choose from $\{imp\_neg_{1},imp\_neg_{2},imp\_neg_{5}\}$. For subsequent distractor templates, we sample from without replacement from these implicit templates.
Appendix C Annotation instructions
C.1 Task 1 Description
@@ C.1.1 Grammaticality
C.1.2 Questions about pronouns
@@ C.2 Task 2 Description
Appendix D Experimental Details
@@
Appendix E Prompting
Table 3 shows all 10 prompt templates we use with chat models. The first five are open-ended and the last five provide the model with pronoun options to choose from. Table 4 shows instantiated examples of one of the templates for both FLAN-T5 and Llama-2-chat. The number of maximum new tokens is set to 5 for FLAN-T5 and 20 for Llama-2-chat based on experiments.
@@ Appendix E Prompting
The technician told the customer that ___ could pay with cash.
@@
The technician told the customer that ___ could pay with cash.
@@
Appendix F Model correlations between pronoun predictions by case
Table 5 shows correlations between pronoun predictions across cases for individual models. The distribution of these scores are visualized in Figure 2.
Model Nom-Acc Nom-Poss Acc-Poss
BERT-base 0.32 0.21 0.31
BERT-large -0.05 0.35 0.03
RoBERTa-base -0.46 0.24 -0.27
RoBERTa-large 0.26 0.37 0.23
ALBERT-base-v2 0.19 -0.16 0.27
ALBERT-large-v2 0.06 0.07 0.22
ALBERT-xlarge-v2 0.26 0.30 0.03
ALBERT-xxlarge-v2 0.41 0.28 0.18
MosaicBERT -0.23 0.21 -0.32
OPT-125M -0.03 0.42 0.03
OPT-350M -0.04 0.35 0.04
OPT-1.3B -0.25 0.39 -0.18
OPT-2.7B -0.12 0.15 -0.27
OPT-6.7B 0.00 0.20 -0.10
OPT-13B -0.08 0.48 -0.07
OPT-30B -0.32 0.23 -0.01
OPT-66B 0.17 0.29 0.17
Pythia-14M 0.12 0.23 0.10
Pythia-70M 0.32 0.29 0.05
Pythia-160M 0.47 0.40 0.03
Pythia-410M 0.41 0.30 0.05
Pythia-1B -0.12 0.49 -0.14
Pythia-1.4B 0.15 0.35 0.11
Pythia-2.8B -0.16 0.42 -0.14
Pythia-6.9B -0.12 0.53 0.06
Pythia-12B 0.02 0.54 -0.07
Llama-2-7B 0.09 0.34 -0.21
llama-2-13B 0.00 0.55 -0.01
llama-2-70B 0.36 0.47 0.02
Table 5: Individual correlations between pronoun predictions across different cases.

### v2

@@ Acknowledgements
The authors thank Timm Dill for several rounds of patient annotation, as well as Aaron Mueller, Marius Mosbach, Vlad Niculae and Yanai Elazar, for feedback and discussions about math, plots and framing. Vagrant Gautam received funding from the BMBF’s (German Federal Ministry of Education and Research) SLIK project under the grant 01IS22015C. The work of Anne Lauscher is funded under the Excellence Strategy of the German Federal Government and States.
Appendix A Limitations
Shallow heuristics. Much of the recent progress on reasoning datasets has been critically investigated and shown to often be a result of spurious correlations and dataset artifacts (Trichelair et al., 2019; Elazar et al., 2021). We caution readers that our dataset also gives a very generous estimate of model reasoning performance for multiple reasons. Many of our task sentences are not “Google-proof” (Levesque et al., 2012), i.e., they can be solved with shallow heuristics such as word co-occurrences. Consider the following task sentence: The janitor said not to step on the wet floor, otherwise ___ would have to mop it all over again. The association between janitor and mop is strong compared to child (the participant paired with janitor) and mop, and could easily be exploited by models to solve the dataset without solving the task with something resembling “reasoning.” Another shallow heuristic that can be used to solve our current dataset is to simply return the first pronoun in the context, which happens to always be the correct answer. Our dataset design is flexible and allows for the creation of other orderings of sentences, but this is another example of why our dataset in its current form should only be used as an evaluation dataset, and models should not be pre-trained or fine-tuned with any splits of our data, nor provided with examples for in-context learning.
Whose bias? Our task as it is defined in Section 2 is much broader than the scope of our dataset. We focus on occupations due to the wide attention they have received in prior literature, but we continue a long tradition of ignoring biases relating to the participants, e.g., child, taxpayer, etc.
Data contamination. We take steps to prevent data contamination following Jacovi et al. (2023), including not releasing our data in plain text, and not evaluating with models behind closed APIs that do not guarantee that our data will not be used to train future models. However, as we cannot guarantee a complete absence of data leakage unless we never release the dataset, we encourage caution in interpreting results on RUFF with models trained on data after March 2024.
Appendix B List of Occupations
@@
Appendix C Context Template Construction
Semantically bleached templates such as He is an accountant are well-established in the literature for testing word embedding associations (Caliskan et al., 2017), but it is unnatural to use more than one consecutively, which is necessary to evaluate robustness in this setting. On the other hand, natural corpora like Levy et al. (2021) contain occupation-specific sentences that cannot be used more generally for a controlled setting across occupations and participants. As a compromise between a controlled setting and naturalistic data, we create context templates with generic themes that apply to all humans, e.g., universal emotions and sensations (hungry/full, tired/energetic, unhappy/happy, etc.).
For each case, we create 10 explicit templates which explicitly demonstrate the coreference between an individual and a pronoun using a subordinate clause, and 10 implicit templates, simple sentences which only contain a pronoun as the subject. The first two sentences of the context (i.e., the introduction and the first distractor) are always sampled from the explicit templates, and the rest are sampled from the implicit templates, as this reflects natural and coherent use of pronouns in discourse.
@@ Appendix C Context Template Construction
After making a choice for the first distractor template, we fill it with any of the three remaining pronouns and then we remove this template’s index from our pool, but re-add the index of the introductory template. This is because subsequent distractor templates always use implicit templates. For example, if we chose $exp\_neg_{4}$ as our first distractor template, we would now choose from $\{imp\_neg_{1},imp\_neg_{2},imp\_neg_{5}\}$. For subsequent distractor templates, we sample without replacement from these implicit templates.
Appendix D Annotator Demographics
All three annotators (two authors and an additional annotator) are fluent English speakers. The two authors who create and validate templates have linguistic training at the undergraduate level. One author and one annotator have experience with using singular they and neopronouns, while the other author has prior exposure to singular they but not the neopronoun xe.
Appendix E Annotation Instructions
E.1 Task 1 Description
@@ E.1.1 Grammaticality
E.1.2 Questions about pronouns
@@ E.2 Task 2 Description
Appendix F Experimental Details
@@
Appendix G Prompting
Table 3 shows all 10 prompt templates we use to present our task instances to chat models. The first five are open-ended and the last five provide the model with pronoun options to choose from. These templates are used within vanilla and chain-of-thought prompting paradigms as described below.
@@ Appendix G Prompting
G.1 Vanilla prompting
With FLAN-T5, vanilla prompting only requires instantiating templates with task instances, whereas Llama-2-chat requires special formatting with INST. Instantiated examples of one template are shown for both models in Table 4. The number of maximum new tokens is set to 5 for FLAN-T5 and 20 for Llama-2-chat based on experimentation.
@@ G.1 Vanilla prompting
The accountant had just eaten a big meal so her stomach
was full. The accountant was asked about ___ charges for
preparing tax returns.
@@
The accountant had just eaten a big meal so her stomach
was full. The accountant was asked about ___ charges for
preparing tax returns.
@@
G.2 Chain-of-thought prompting
We focus on FLAN-T5-xxl for chain-of-thought experiments, and use the strong zero-shot prompts for reasoning proposed by Kojima et al. (2022) (“Let’s think step by step”) and Zhou et al. (2023) (“Let’s work this out in a step by step way to be sure we have the right answer”), which we append after the template. Following their codebases, we first allow the models to generate a chain of thought (with 128 maximum new tokens). Then, we append the chain of thought after the question and elicit the final answer with the string “Therefore, the correct pronoun is,” allowing the model to generate up to 10 new tokens. This two-step process is illustrated with examples in Table 5. We save both the final answer and the chain of thought for later analysis.
Chain-of-thought prompt Example input
Kojima et al. (2022) Q: Fill in the blank with the appropriate pronoun.
(eliciting chain of thought)
The accountant had just eaten a big meal so her stomach
was full. The accountant was asked about ___ charges for
preparing tax returns.
A: Let’s think step by step.
Kojima et al. (2022) Q: Fill in the blank with the appropriate pronoun.
(eliciting final answer)
The accountant had just eaten a big meal so her stomach
was full. The accountant was asked about ___ charges for
preparing tax returns.
A: Let’s think step by step. {generated chain of thought}
Therefore, the correct pronoun is
Zhou et al. (2023) Q: Fill in the blank with the appropriate pronoun.
(eliciting chain of thought)
The accountant had just eaten a big meal so her stomach
was full. The accountant was asked about ___ charges for
preparing tax returns.
A: Let’s work this out in a step by step way to be sure we
have the right answer.
Zhou et al. (2023) Q: Fill in the blank with the appropriate pronoun.
(eliciting final answer)
The accountant had just eaten a big meal so her stomach
was full. The accountant was asked about ___ charges for
preparing tax returns.
A: Let’s work this out in a step by step way to be sure we
have the right answer. {generated chain of thought}
Therefore, the correct pronoun is
Table 5: Example input using template 3 for evaluating FLAN-T5-xxl with two types of chain-of-thought prompting. Prompting happens in two phases regardless of the choice of prompt: eliciting the chain of thought and eliciting the final answer.
he/him/his Refer to caption
she/her/her Refer to caption
they/them/their Refer to caption
xe/xem/xyr Refer to caption
Figure 8: Counts of pronoun predictions from all models, in the absence of context. The random baseline shows counts if each pronoun set was chosen equally often.
Appendix H Context-free pronoun predictions by model
Figure 8 shows per-model pronoun predictions in the absence of context. All models predict he/him/his more frequently than she/her/her, which is in turn predicted more frequently than they/them/their and xe/xem/xyr. However, encoder-only models are more balanced in their predictions across the four pronoun sets, compared to decoder-only models which show very stark differences in pronoun predictions.
Appendix I Results with Vanilla and Chain-of-Thought Prompting
I.1 Vanilla prompting
Prompting is a different model evaluation mechanism than log likelihoods, with higher task demands that lead to lower performance than log likelihoods with both base models and instruction fine-tuned chat models (Hu and Levy, 2023; Hu and Frank, 2024; Kauf et al., 2024). We thus expect vanilla prompting results (using the prompts listed in Appendix G) to be worse than results with log likelihoods. Indeed, Figure 9(a) shows that Llama-2-chat prompting performance is lower than Llama-2 evaluated with log likelihoods, even with no distractors. Figure 9(b) shows the results of standard prompting with FLAN-T5, an encoder-decoder model which shows similar patterns of degradation to decoder-only models. Bigger models are mostly better and degrade more gracefully than the smaller ones, but there remains a lot of variance across prompts, as shown in the box plots.
(a) Llama-2-chat
(b) FLAN-T5
Figure 9: Performance of chat models (Llama-2-chat and FLAN-T5) with additional distractors, using vanilla prompting. The boxplots show the range of performance across 10 different templates.
I.2 Chain-of-thought prompting
As FLAN-T5-xxl shows strong performance with low variance compared to all the other chat models we consider, we focus on this model for additional evaluation with chain-of-thought prompting. Zero-shot chain-of-thought prompting encourages models to think step-by-step, which could in theory produce much better results on pronoun fidelity. While chain-of-thought prompting is excessive for a task as simple as pronoun fidelity, it might encourage the model to explicitly list the referents and associated pronouns, which could help the model predict the correct pronoun with higher accuracy. In practice, however, we find that it leads to worse performance, potentially due to hallucination.
Figure 10 shows the pronoun fidelity of FLAN-T5-xxl with different types of prompting based on the final answer the model provides. Both types of chain-of-thought prompting worsen performance and increase the variance across prompts compared to vanilla prompting.
When examining model-generated answers and chains of thought, we found that FLAN-T5-xxl does not in fact solve the problem step by step as the instruction suggests. Instead, the chain of thought often already contains an answer, and the final answer is not necessarily the same as this one. Therefore, we also plot performance using answers from the model-generated chain of thought in Figure 11. Once again, performance with 1-5 distractors is much lower, showing that chain-of-thought prompting degrades performance compared to vanilla prompting. However, with no distractors, performance is almost exactly the same as vanilla prompting, as models simply generate the answer within the chain of thought. This reinforces that chain-of-thought is unnecessary for a task this simple.
Figure 10: Performance of FLAN-T5-xxl with distractor sentences, comparing vanilla prompting to two types of chain-of-thought prompting. Here, the model’s final answers are used for evaluation and the boxplots show the range of performance across 10 different templates.
Figure 11: Performance of FLAN-T5-xxl with distractor sentences, comparing vanilla prompting to two types of chain-of-thought prompting. Here, the model’s chain of thought is used for evaluation and the boxplots show the range of performance across 10 different templates.